Charisma, Inc., has debt outstanding with a face value of $6.4 million. The value of the firm if it were entirely financed by equity would be $31.4 million. The company also has 435,000 shares of stock outstanding that sell at a price of $60 per share. The corporate tax rate is 24 percent. What is the decrease in the value of the company due to expected bankruptcy costs? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567.)
Answer #1

Value of levered company = Value of unlevered equity + tax(debt)

Value of levered company = 31,400,000 + 0.24(6,400,000)

Value of levered company = 31,400,000 + 1,536,000

Value of levered company = 32,936,000

Market value of firm = (435,000 * 60) + 6,400,000 = 32,500,000

Decrease = 32,936,000 - 32,500,000

Decrease = $436,000